# Trial by FIRE: Probing the dark matter density profile of dwarf galaxies with GraphNPE

### v1

@@ Abstract
The Dark Matter (DM) distribution in dwarf galaxies provides crucial insights into both structure formation and the particle nature of DM. GraphNPE (Graph Neural Posterior Estimator), first introduced in Nguyen et al. (2023), is a novel simulation-based inference framework that combines graph neural networks and normalizing flows to infer the DM density profile from line-of-sight stellar velocities. Here, we apply GraphNPE to satellite dwarf galaxies in the FIRE-2 Latte simulation suite of Milky Way-mass halos, testing it against both Cold and Self-Interacting DM scenarios. Our method demonstrates superior precision compared to conventional Jeans-based approaches, recovering DM density profiles to within the 95% confidence level even in systems with as few as 30 tracers. Moreover, we present the first evaluation of mass modeling methods in constraining two key parameters from realistic simulations: the peak circular velocity, $V_{\mathrm{max}}$, and the peak virial mass, $M_{\mathrm{200m}}^{\mathrm{peak}}$. Using only line-of-sight velocities, GraphNPE can reliably recover both $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ within our quoted uncertainties, including those experiencing tidal effects ($\gtrsim 63\%$ of systems are recovered with our 68% confidence intervals and $\gtrsim 92\%$ within our $95\%$ confidence intervals). The method achieves $10-20\%$ accuracy in $V_{\mathrm{max}}$ recovery, while $M_{\mathrm{200m}}^{\mathrm{peak}}$ is recovered to $0.1-0.4\,\mathrm{dex}$ accuracy. This work establishes GraphNPE as a robust tool for inferring DM density profiles in dwarf galaxies, offering promising avenues for constraining DM models. The framework’s potential extends beyond this study, as it can be adapted to non-spherical and disequilibrium models, showcasing the broader utility of simulation-based inference and graph-based learning in astrophysics.
@@
pubyear: 2025pagerange: Trial by FIRE: Probing the dark matter density profile of dwarf galaxies with GraphNPEC.2
@@ 1 Introduction
Dark matter (DM) makes up about 85% of matter in the Universe, playing an integral role in the formation and evolution of galaxies and clusters (White & Rees, 1978; Navarro et al., 1997; Subramanian et al., 2000; Dekel et al., 2003; Springel et al., 2005). Despite its significance, the particle nature of DM remains among the greatest outstanding questions in modern astrophysics and cosmology. While the standard Lambda Cold Dark Matter ($\Lambda$CDM) paradigm has successfully explained the large-scale structure of the Universe, such as the fluctuations in the cosmic microwave background (CMB)  (e.g. Bennett et al., 2013; Planck Collaboration et al., 2020) and the distribution of galaxy clusters (e.g. Tinker et al., 2008; Rozo et al., 2010), discrepancies arise at the small scales. Observations of dwarf galaxies in the Local Group have uncovered small-scale challenges, including the core-cusp problem (Flores & Primack, 1994; Moore, 1994; Navarro et al., 1996; Read & Gilmore, 2005; Spekkens et al., 2005; Oh et al., 2011; Pontzen & Governato, 2014; Oh et al., 2015; Read et al., 2019; De Leo et al., 2024), missing satellites problem (Klypin et al., 1999; Moore et al., 1999; Read & Erkal, 2019), too-big-to-fail (Boylan-Kolchin et al., 2011), satellite plane (Kunkel & Demers, 1976; Lynden-Bell, 1976), and diversity of shapes problem (Oman et al., 2015; Creasey et al., 2017; Zavala et al., 2019; Hayashi et al., 2020a) (for a review, refer to Bullock & Boylan-Kolchin, 2017). These challenges can be partially addressed by a more complete understanding of baryonic processes (Governato et al., 2010; Pontzen & Governato, 2012; Fitts et al., 2017; Garrison-Kimmel et al., 2017; Kim et al., 2018; Simpson et al., 2018; Garrison-Kimmel et al., 2019; Buck et al., 2019; Sales et al., 2022), environmental effects such as tidal interaction with the Milky Way and other galaxies (Boylan-Kolchin et al., 2011; Mayer et al., 2006; Read et al., 2006), and survey selection effects (Tollerud et al., 2008; Drlica-Wagner et al., 2020). However, open questions still remain (Pawlowski et al., 2015; Santos-Santos et al., 2020), motivating alternative theories of DM such as self-interaction (Spergel & Steinhardt, 2000).
Characterized by their high mass-to-light ratio, small size, and proximity to the Galaxy, the Milky Way’s satellite galaxies present an exciting frontier for exploring and testing the particle nature of DM. Key observables, including the satellite’s inner DM density profile, their mass function, luminosity function, and phase-space distribution, can offer stringent constraints on DM properties, such as self-interaction cross-section (Rocha et al., 2013; Kaplinghat et al., 2016; Tulin & Yu, 2018; Read et al., 2018; Correa, 2021; Toguz et al., 2022; Júlio et al., 2023), DM dissipation (Kaplan et al., 2010; Cyr-Racine & Sigurdson, 2013; Shen et al., 2021, 2024; Gemmell et al., 2024; Roy et al., 2024), DM-baryon interaction (Nadler et al., 2019), and mass of warm and fuzzy DM (Bode et al., 2001; Kennedy et al., 2014; Lovell et al., 2014; Bose et al., 2016; Alvey et al., 2021; Hayashi et al., 2021; Dalal & Kravtsov, 2022; Tan et al., 2024).
@@
Baryonic processes, such as feedback, can also significantly influence the density profiles (e.g. Di Cintio et al., 2014a, b; Read et al., 2019; De Leo et al., 2024). Feedback can lead to a diverse range of cored and cuspy profiles, depending on where a galaxy is in its cycle of gas inflows and outflows (El-Badry et al., 2016, 2017; Muni et al., 2025). However, if the gas is stripped, such as in Milky Way satellites, the galaxy can be “frozen into” a cored state. This effect is particularly relevant for a specific mass range, including classical dwarfs and Milky Way-mass galaxies (e.g. Lazar et al., 2020). Consequently, dwarf galaxies, particularly the ultra-faint, are the ideal laboratory for investigating the inner profile affected solely by the particle nature of DM (e.g. Zoutendijk et al., 2021b, a).
@@
Apart from serving as a direct probe of the particle nature of DM through its inner slope of the DM density profile, dwarf galaxies are an excellent target for indirect detection of DM. They are relatively close, making the annihilation/decay rates of DM into standard model particles quite high and more easily detectable, and generally devoid of astrophysical backgrounds that might mimic DM signals (e.g., the potential population of millisecond pulsars near the Galactic Center muddling the waters as to whether DM is the source of the Galactic center excess (Cirelli et al., 2011; Ackermann et al., 2011; Geringer-Sameth & Koushiappas, 2011; Mazziotta et al., 2012; Ackermann et al., 2014; Martinez, 2015; Ackermann et al., 2015; Geringer-Sameth et al., 2015a, b; Anderson et al., 2015; Leane & Slatyer, 2019; Ajello et al., 2024)).
@@
Given the critical information that is encoded in the density profile of DM, one needs to work through recovering such profiles, particularly in dwarf galaxies where the effect of baryons is minimal. To tackle such problem, one could recover the underlying gravitational potential of the galaxy through the dynamics of its stars. Traditionally, this has been done using the Jeans dynamical modeling techniques (Jeans, 1915; Bonnivard et al., 2015), which relate the velocity dispersion of tracer stars within these galaxies to their DM-dominated gravitational potential. Jeans methods are straightforward to implement. Most models assume spherical symmetry, but tests on non-spherical mocks show that for less than $\sim 10,000$ tracer stars, this is reasonable approximation(Read & Steger, 2017; Genina et al., 2020). Some models have explored axisymmetric fits to nearby dwarfs (Hayashi et al., 2020b, 2023), while even highly tidally disrupting dwarfs have been shown to be successfully modeled, providing contaminating tidal debris can be correctly distinguished from bound stars (Read et al., 2018; De Leo et al., 2024).
@@
A long-standing challenge for Jeans methods, applied to just line-of-sight velocity data, has been mass-anisotropy degeneracy (Merrifield & Kent, 1990; Wilkinson et al., 2002; Łokas & Mamon, 2003; de Lorenzi et al., 2009; Read & Steger, 2017; El-Badry et al., 2017; Genina et al., 2020). This refers to a degeneracy between the enclosed mass, that we want to infer, $M(<r)$, and the velocity anisotropy parameter $\beta(r)$ ($\beta=0,1,-\infty$ corresponds to an isotropic, radial anisotropy, and tangential anisotropy velocity dispersion, respectively; Binney & Tremaine 2008). However, this degeneracy can be broken in a number of ways. Firstly, where available, proper motion data can be used (Strigari et al., 2007b; Read & Steger, 2017; Strigari et al., 2018; Massari et al., 2020; Vitral et al., 2024). Secondly, we can use higher order moments of the velocity distribution, for example through “Virial Shape Parameters” (Merrifield & Kent, 1990; Łokas, 2009; Richardson & Fairbairn, 2014; Read & Steger, 2017; Mamon et al., 2013). Thirdly, we can simultaneously fit multiple tracer populations moving within the same gravitational potential (e.g. Walker & Peñarrubia, 2011; Amorisco & Evans, 2012; Zhu et al., 2016; Read & Steger, 2017). And finally, we can move away from Jeans modeling to Schwarzschild (e.g. Breddels & Helmi, 2013, 2014; Kowalczyk et al., 2019) or distribution function modeling (Wilkinson et al., 2002; Pascale et al., 2018; Read et al., 2021). These latter methods simultaneously fit all moments of the velocity distribution function, $f(v)$, by assuming some functional form for $f$ (or a range of functional forms).
@@
We also investigate the recovery of the peak circular velocity of halos, $V_{\mathrm{max}}$, and their peak virial mass, $M_{\mathrm{200m}}$–both of key interest for constraining cosmological models (e.g. Burkert, 1995; Bullock et al., 2001; van den Bosch et al., 2003; Rodríguez-Puebla et al., 2016; Read & Erkal, 2019; Kim & Peter, 2021). This is the first study to evaluate how well mass modeling methods can constrain $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$ for simulated dwarf galaxies in realistic environments. This test is particularly timely given that two state-of-the-art mass modeling methods in the literature, CJAM (Jeans Anisotropic Multi-Gaussian Expansion; Watkins et al. 2013) and GravSphere, yield quite different inferences when applied to the same data (Zoutendijk et al., 2021a).
@@ 2.1 FIRE-2 Simulations
In this paper, we aim to test the validity of GraphNPE on a more realistic scenarios than previously done in N23, with the end goal of applying GraphNPE to the Milky Way’s satellites. To that end, we use the satellite population of the Milky Way-mass simulated galaxies as part of the Latte suite (Wetzel et al., 2016). The Latte suite is a set of zoom-in simulations of Milky Way-mass galaxies, in which the hydrodynamics is based on the FIRE-2 physics (Hopkins et al., 2018). In order to ensure robustness of the results in a varied density profiles, we will use the CDM galaxies (Wetzel et al., 2016, 2023), along with SIDM galaxies presented in (Sameie et al., 2021; Vargya et al., 2022; Arora et al., 2024). We discuss each of these properties in turn.
FIRE-2 physics: The simulations are run using the GIZMO [^3] https://bitbucket.org/phopkins/gizmo-public code base (Hopkins, 2014, 2015), which utilizes a mesh-free, finite-mass (MFM) hydrodynamic solver, along with a version of the Tree-PM gravity solver from GADGET-3 (Springel, 2005). MFM is a mesh-free Lagrangian finite-volume Godunov method that provides adaptive spatial resolution while maintaining conservation of mass, energy, momentum, and angular momentum. It integrates the advantages of both smoothed particle hydrodynamics (SPH) and traditional grid-based methods, enabling adaptive resolution and the accurate capture of fluid dynamics in highly complex astrophysical phenomena. The FIRE-2 physics model (Hopkins et al., 2018) additionally includes star formation (Hopkins et al., 2013), stellar feedback from radiation pressure, supernovae Ia and II, stellar winds, photoelectric heating, and photoionization, along with a detailed treatment of cooling, heating, UV background (Faucher-Giguère et al., 2009), metal mixing in the interstellar medium (Su et al., 2017; Escala et al., 2018) and enrichment from stars and supernovae (Ma et al., 2016).
@@
Dark Matter Models: To make sure that we span a variety of inner profiles, we include in our analysis simulations with Cold Dark Matter (CDM) as well as Self-Interacting Dark Matter (SIDM) (see e.g. Tulin & Yu, 2018, for a review). We analyze the fiducial CDM simulations from the FIRE-2 public data release (Wetzel et al., 2023) and the SIDM simulations presented in Sameie et al. (2021); Vargya et al. (2022); Arora et al. (2024). [^4] Prior SIDM simulations using the FIRE-2 galaxy formation model include simulations of $10^{10}\,\mathrm{M_{\odot}}$ halos (Robles et al., 2017; Fitts et al., 2019) and dissipative SIDM scenarios (Shen et al., 2021, 2024); however, these are not included in this study. For the SIDM simulations, we use two hosts, m12f and m12m, each simulated with two constant DM self-interaction cross-sections of $\sigma/m=1$ and $10\,\mathrm{cm^{2}/g}$. DM self-interaction is implemented using the Monte Carlo scattering algorithm in Rocha et al. (2013); Peter et al. (2013), which determines the probability of interaction using a spline kernel with adaptive smoothing length (Monaghan & Lattanzio, 1985). All scattering events are modeled as elastic and isotropic in the center of mass frame.
@@
$[\mathrm{M_{\odot}}]$ $[\mathrm{kpc}]$ $[\mathrm{cm^{2}/g}]$

### v2

@@ Abstract
The Dark Matter (DM) distribution in dwarf galaxies provides crucial insights into both structure formation and the particle nature of DM. GraphNPE (Graph Neural Posterior Estimator), first introduced in Nguyen et al. (2023), is a novel simulation-based inference framework that combines graph neural networks and normalizing flows to infer the DM density profile from line-of-sight stellar velocities. Here, we apply GraphNPE to satellite dwarf galaxies in the FIRE-2 Latte simulation suite of Milky Way-mass halos, testing it against both Cold and Self-Interacting DM scenarios. Our method demonstrates superior precision compared to conventional Jeans-based approaches, recovering DM density profiles to within the 95% confidence level even in systems with as few as 30 tracers. Moreover, we present the first evaluation of mass modeling methods in constraining two key parameters from realistic simulations: the peak circular velocity, $V_{\mathrm{max}}$, and the peak virial mass, $M_{\mathrm{200m}}^{\mathrm{peak}}$. Using only line-of-sight velocities, GraphNPE can reliably recover both $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ within our quoted uncertainties, including those experiencing tidal effects ($\gtrsim 63\%$ of systems are recovered within our 68% confidence intervals and $\gtrsim 92\%$ within our $95\%$ confidence intervals). The method achieves $10-20\%$ accuracy in $V_{\mathrm{max}}$ recovery, while $M_{\mathrm{200m}}^{\mathrm{peak}}$ is recovered to $0.1-0.4\,\mathrm{dex}$ accuracy. This work establishes GraphNPE as a robust tool for inferring DM density profiles in dwarf galaxies, offering promising avenues for constraining DM models. The framework’s potential extends beyond this study, as it can be adapted to non-spherical and disequilibrium models, showcasing the broader utility of simulation-based inference and graph-based learning in astrophysics.
@@
pubyear: 2025pagerange: Trial by FIRE: Probing the dark matter density profile of dwarf galaxies with GraphNPEC.3
@@ 1 Introduction
[…]
Characterized by their high mass-to-light ratio, small size, and proximity to the Galaxy, the Milky Way’s satellite galaxies present an exciting frontier for exploring and testing the particle nature of DM. Key observables, including the satellite’s inner DM density profile, their mass function, luminosity function, and phase-space distribution, can offer stringent constraints on DM properties, such as self-interaction cross-section (Rocha et al., 2013; Kaplinghat et al., 2016; Tulin & Yu, 2018; Read et al., 2018; Correa, 2021; Toguz et al., 2022; Júlio et al., 2023), DM dissipation (Kaplan et al., 2010; Cyr-Racine & Sigurdson, 2013; Foot & Vagnozzi, 2015b, a, 2016; Shen et al., 2021, 2024; Gemmell et al., 2024; Roy et al., 2024), DM-baryon interaction (Nadler et al., 2019), and mass of warm and fuzzy DM (Bode et al., 2001; Kennedy et al., 2014; Lovell et al., 2014; Bose et al., 2016; Alvey et al., 2021; Hayashi et al., 2021; Dalal & Kravtsov, 2022; Tan et al., 2024).
@@
Baryonic processes, such as feedback, can also significantly influence the density profiles (e.g. Di Cintio et al., 2014a, b; Read et al., 2019; De Leo et al., 2024). Feedback can lead to a diverse range of cored and cuspy profiles, depending on where a galaxy is in its cycle of gas inflows and outflows (El-Badry et al., 2016, 2017; Muni et al., 2025). However, if the gas is stripped, such as in Milky Way satellites, the galaxy can be “frozen into” a cored state. This effect is particularly relevant for a specific mass range, including classical dwarfs and Milky Way-mass galaxies (e.g. Lazar et al., 2020). Consequently, dwarf galaxies are the ideal laboratory for investigating the inner profile affected solely by the particle nature of DM (e.g. Zoutendijk et al., 2021b, a). This is particularly true for ultra-faint dwarfs ($M_{V}\gtrsim-7.7$), which experience weaker stellar feedback.
@@
Apart from serving as a direct probe of the particle nature of DM through its inner slope of the DM density profile, Dwarf galaxies are an excellent target for indirect detection of DM. They are relatively close, making the annihilation/decay rates of DM into standard model particles quite high and more easily detectable, and generally devoid of astrophysical backgrounds that might mimic DM signals (e.g., the potential population of millisecond pulsars near the Galactic Center muddling the waters as to whether DM is the source of the Galactic center excess (Cirelli et al., 2011; Ackermann et al., 2011; Geringer-Sameth & Koushiappas, 2011; Mazziotta et al., 2012; Ackermann et al., 2014; Martinez, 2015; Ackermann et al., 2015; Geringer-Sameth et al., 2015a, b; Anderson et al., 2015; Leane & Slatyer, 2019; Ajello et al., 2024).
@@
Given the critical information that is encoded in the density profile of DM, one needs to work through recovering such profiles, particularly in dwarf galaxies where the effect of baryons is minimal. To tackle this problem, one could recover the underlying gravitational potential of the galaxy through the dynamics of its stars. Traditionally, this has been done using the Jeans dynamical modeling techniques (Jeans, 1915; Bonnivard et al., 2015), which relate the velocity dispersion of tracer stars within these galaxies to their DM-dominated gravitational potential. Jeans methods are straightforward to implement. Most models assume spherical symmetry, which has been demonstrated to be a reasonable approximation on non-spherical mocks with less than $\sim 10,000$ tracer stars (Read & Steger, 2017; Genina et al., 2020). Some models have explored axisymmetric fits to nearby dwarfs (Hayashi et al., 2020b, 2023). Even highly tidally disrupting dwarfs have been shown to be successfully modeled, providing contaminating tidal debris can be correctly distinguished from bound stars (Read et al., 2018; De Leo et al., 2024).
@@
A long-standing challenge for Jeans methods, applied to just line-of-sight velocity data, has been mass-anisotropy degeneracy (Merrifield & Kent, 1990; Wilkinson et al., 2002; Łokas & Mamon, 2003; de Lorenzi et al., 2009; Read & Steger, 2017; El-Badry et al., 2017; Genina et al., 2020). This refers to a degeneracy between the enclosed mass, that we want to infer, $M(<r)$, and the velocity anisotropy parameter $\beta(r)$ ($\beta=0,1,$-$\infty$ correspond to an isotropic, radial anisotropy, and tangential anisotropy velocity dispersion, respectively; Binney & Tremaine 2008). However, this degeneracy can be broken in a number of ways. Firstly, where available, proper motion data can be used (Strigari et al., 2007b; Read & Steger, 2017; Strigari et al., 2018; Massari et al., 2020; Vitral et al., 2024). Secondly, we can use higher order moments of the velocity distribution, for example through “Virial Shape Parameters” (Merrifield & Kent, 1990; Łokas, 2009; Richardson & Fairbairn, 2014; Read & Steger, 2017; Mamon et al., 2013). Thirdly, we can simultaneously fit multiple tracer populations moving within the same gravitational potential (e.g. Walker & Peñarrubia, 2011; Amorisco & Evans, 2012; Zhu et al., 2016; Read & Steger, 2017). And finally, we can move away from Jeans modeling to Schwarzschild (e.g. Breddels & Helmi, 2013, 2014; Kowalczyk et al., 2019) or distribution function modeling (Wilkinson et al., 2002; Pascale et al., 2018; Read et al., 2021). These latter methods simultaneously fit all moments of the velocity distribution function, $f(v)$, by assuming some functional form for $f$ (or a range of functional forms).
@@
We also investigate the recovery of the peak circular velocity of halos, $V_{\mathrm{max}}$, and their peak virial mass, $M_{\mathrm{200m}}^{\mathrm{peak}}$–both of key interest for constraining cosmological models (e.g. Burkert, 1995; Bullock et al., 2001; van den Bosch et al., 2003; Rodríguez-Puebla et al., 2016; Read & Erkal, 2019; Kim & Peter, 2021). This is the first study to evaluate how well mass modeling methods can constrain $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ for simulated dwarf galaxies in realistic environments. This test is particularly timely given that two state-of-the-art mass modeling methods in the literature, CJAM (Jeans Anisotropic Multi-Gaussian Expansion; Watkins et al. 2013) and GravSphere, yield quite different inferences when applied to the same data (Zoutendijk et al., 2021a).
@@ 2.1 FIRE-2 Simulations
In this paper, we aim to test the validity of GraphNPE on more realistic scenarios than previously done in N23, with the end goal of applying GraphNPE to the Milky Way’s satellites. To that end, we use the satellite population of the Milky Way-mass simulated galaxies as part of the Latte suite (Wetzel et al., 2016). The Latte suite is a set of zoom-in simulations of Milky Way-mass galaxies, in which the hydrodynamics is based on the FIRE-2 physics (Hopkins et al., 2018). In order to ensure robustness of the results with varying density profiles, we will use the CDM galaxies (Wetzel et al., 2016, 2023), along with SIDM galaxies presented in (Sameie et al., 2021; Vargya et al., 2022; Arora et al., 2024). We discuss each of these properties in turn.
FIRE-2 physics: The simulations are run using the GIZMO [^3] https://bitbucket.org/phopkins/gizmo-public code base (Hopkins, 2014, 2015), which utilizes a mesh-free, finite-mass (MFM) hydrodynamic solver, along with a version of the Tree-PM gravity solver from GADGET-3 (Springel, 2005). MFM is a mesh-free Lagrangian finite-volume Godunov method that provides adaptive spatial resolution while maintaining conservations of mass, energy, momentum, and angular momentum. It integrates the advantages of both smoothed particle hydrodynamics (SPH) and traditional grid-based methods, enabling adaptive resolution and the accurate capture of fluid dynamics in highly complex astrophysical phenomena. The FIRE-2 physics model (Hopkins et al., 2018) additionally includes star formation (Hopkins et al., 2013), stellar feedback from radiation pressure, supernovae Ia and II, stellar winds, photoelectric heating, and photoionization, along with a detailed treatment of cooling, heating, UV background (Faucher-Giguère et al., 2009), metal mixing in the interstellar medium (Su et al., 2017; Escala et al., 2018) and enrichment from stars and supernovae (Ma et al., 2016).
@@
Dark Matter Models: To make sure that we span a variety of inner profiles, we include in our analysis simulations with Cold Dark Matter (CDM) as well as Self-Interacting Dark Matter (SIDM) (see e.g. Tulin & Yu, 2018, for a review). We analyze the fiducial CDM simulations from the FIRE-2 public data release (Wetzel et al., 2023) and the SIDM simulations presented in Sameie et al. (2021); Vargya et al. (2022); Arora et al. (2024). [^4] Prior SIDM simulations using the FIRE-2 galaxy formation model include simulations of $10^{10}\,\mathrm{M_{\odot}}$ halos (Robles et al., 2017; Fitts et al., 2019) and dissipative SIDM scenarios (Shen et al., 2021, 2024); however, these are not included in this study. For the SIDM simulations, we use two hosts, m12f and m12m, each simulated with two constant DM self-interaction cross-sections of $\sigma/m=1$ and $10\,\mathrm{cm^{2}/g}$. DM self-interaction is implemented using the Monte Carlo scattering algorithm in Rocha et al. (2013); Peter et al. (2013), which determines the probability of interaction using a spline kernel with adaptive smoothing length (Monaghan & Lattanzio, 1985). All scattering events are modeled as elastic and isotropic in the center-of-mass frame.
@@
$\mathrm{M_{\odot}}] [$$\mathrm{kpc}] [$$\mathrm{cm^{2}/g}]$